A current of 6A is flowing at 220V in the primary coil of a transformer.If the voltage produced in the secondary coil is 1100V and 40% of power is lost then the current in the secondary coil will be:
1
Expert's answer
2020-10-08T18:14:38-0400

Explanations & Calculations


  • As 40% power is lost during the usage, the available power for the secondary or the efficiency of the transformer is 60%.
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small 220V\times 6A \times \frac{60}{100}&= \small 1100V\times i_s\\
i_s &= \small \bold{0.72A = 720mA }
\end{aligned}"